# Optimal parallelisation strategies for flat histogram Monte Carlo sampling

### v2

@@ II.3 Model Implementation
As a test case for the parallelisation strategies studied in this work, we consider the application of WL sampling to the study of the phase behaviour of multicomponent, high-entropy alloys (HEAs). The application of WL to multicomponent alloys is an emerging area of research [48, 49, 50] and is therefore a relevant test case for this study.
@@
In this work, we consider the application of WL sampling to a fixed lattice model of the AlTiCrMo high-entropy alloy as a benchmark system, which is classified as a refractory high-entropy superalloy [57] and forms a (partially) chemically ordered structure with decreasing temperature [58, 59]. The parallelisation schemes considered in this study are implemented within the BraWl package [60], which focuses specifically on fixed-lattice simulations of substitutional alloys. In particular, the code implements the Bragg–Williams model [61, 62], which describes the internal energy of a given alloy as a sum over atom-atom effective pair interactions, which are assumed to have finite range. Explicitly, this Hamiltonian takes the form
@@
where $V_{i\alpha;j\alpha^{\prime}}$ denotes the effective pair interaction (EPI) between an atom of chemical species $\alpha$ on lattice site $i$ and an atom of chemical species $\alpha^{\prime}$ on lattice site $j$. $\{\xi_{i\alpha}\}$ indicates the site occupancies, where $\xi_{i\alpha}=1$ if site $i$ is occupied by an atom of chemical species $\alpha$, and $\xi_{i\alpha}=0$ otherwise. It is helpful to observe that this Hamiltonian can be thought of as a generalisation of the Lenz–Ising model used in elementary studies of magnetic phase transitions [63]. The EPIs of Eq. 12 are generally assumed to be homogeneous, isotropic, and of finite range, which simplifies evaluation of this expression. In this work, the EPIs used are those obtained in Ref. [59] for the AlTiCrMo HEA using the $S^{(2)}$ theory for multicomponent alloys [64, 65, 66].
@@ III Results and Discussion
In order to investigate the various outlined parallelisation schemes, the possible parallelisation schemes are grouped and labelled as ‘methods’, as shown in Table 1. Once grouped, the effect of the schemes on speedup per WL instance, efficiency per sub-domain and efficiency per walker, was investigated over a range of $h$ and $m$. The specific model system being simulated was equiatomic AlTiCrMo HEA, the phase behaviour of which has been investigated in an earlier work  [59]. In this benchmarking study assessing the efficiency of various parallelisation schemes, we use a simulation supercell made up of $4\times 4\times 4$ bcc cubic unit cells for a total of 128 atoms. We find that 128 atoms is sufficient to represent the sampling challenge with tests at larger lattice sizes ($6\times 6\times 6$ bcc cubic unit cells, 432 atoms) confirming the trends illustrated here. Our WL sampling simulations used $N=512$ energy bins over an energy domain of E=99E=-99 meV/atom to $E=0$ meV/atom, completing 12 full WL iterations starting with $\log f_{1}=0.05$, where each iteration used half of the logarithm of the modification factor of the previous iteration. The AlTiCrMo HEA model is well-suited as a benchmark system as it exhibits two phase transitions which occur at two separate temperatures [59], and hence convergence of the DOS is non-trivial. We investigated performance for a range of parameters within our parallelisation schemes with the number of walkers per sub-domain $m$ varying from 1 to 6, the number of sub-domains $h$ varying from 1 to 16 and the sub-domain overlap varying from 0% to 75% for both overlap schemes. It is important to note that we impose a requirement that a minimum of two bins must lie in the overlap region. As such, the case which we refer to as 0% overlap still has two bins in the overlap region. This minimum overlap does not change based on the sub-domain size.
@@
III.1 Effect of energy sub-domain overlap
Figure 2: Comparison of the speedup per WL instance for different choices of energy sub-domain overlap, taken relative to the case of a single walker across the entire energy domain. This plot is for simulations using Method 3 with a single walker per energy sub-domain. The dotted red line shows the square of the number of energy sub-domains, i.e., the number of WL instances used squared ($h^{2}$), while the solid red line shows the number of energy sub-domains, i.e., the number of WL instances ($h$). It can be seen that, for all choices of energy sub-domain overlap region size (excluding the 75% case), the speedup is significantly above 100% and that the choice of size of overlap region between energy sub-domains has noticeable impact on speedup.
@@ III.1 Effect of energy sub-domain overlap
Of the overlaps, the 25% to 50% overlap cases appears to be the ones with the most speedup, although not by a statistically significant amount. Such sub-domain overlap allows for better combining of the DOS, as well as a larger region for replica exchange, while still allowing walkers to focus more narrowly on their assigned sub-domain, resulting in more efficient sampling and higher speedup. The other cases introduce either too much or too little overlap, with too little overlap potentially harming the DOS combining and too mush overlap leading to a greater portion of the configuration space being sampled by multiple walkers, which reduces the benefits of parallelisation. For the remainder of this study a 25% overlap is used. This deliberate choice allows for evaluation of performance with an overlap width which remains optimal to within the statistical uncertainty of our data, but still provides a significant range of overlap to realise any benefit from including replica exchange.
@@
It is important to note the super-linear scaling observed in some cases. This greater than 100% speedup per WL instance arises from the reduced number of total MC steps required when sampling due to the energy domain decomposition as explained in the previous section. This effect was discussed above, in Sec. II.4, and is also examined further in Sec. III.4, below. Of further note is the speedup for the first 3 WL instances matches the theoretical limit of $h^{2}$, derived in Sec. II.4 to within the uncertainty of our measurements.
@@
Figure 3: Normalised root mean square error (RMSE) as a function of the percentage overlap. The RMSE was obtained by comparing 23 iteration sampling runs with 16 windows to a well-converged 36 iteration sampling run with 1 window. The error bars correspond to the standard deviation of time taken of 5 repeat WL simulations. For each number of overlap bins, 5 separate 23 iteration simulations were performed and averaged. These data show that there is no correlation between the size of the energy sub-domain overlap region and the accuracy of the obtained global DOS up to a 50% overlap. Beyond 50% overlap the averaging power of the sampling is reduced hence the increase in normalised RMSE though not to a statistically significant extent.
@@
To assess the effect of overlap on DOS convergence, in Figure 3 we show the normalised root mean square error (RMSE) between a 36 iteration (well-converged) WL simulation in one sub-domain and 23 iteration WL simulations of varying overlap size with 16 sub-domains. It can be seen that there is no correlation between DOS convergence accuracy and the percentage overlap in the energy sub-domain overlap region up to 50%. Beyond this point the averaging power of the overlap regions is decreased resulting in an increased normalised RMSE. Of note, however, is that this increase in normalised RMSE increases from 0.15% to 0.3% which is not a statistically significant increase. While remaining at or below a 50% overlap, that the primary benefit of the energy sub-domain overlap region(s) is to facilitate replica exchange and does not harm the sampling.
III.2 Influence of number of walkers used per energy sub-domain
@@
Figure 4: Comparison of the speedup per WL instance for different numbers of walkers per energy sub-domain, taken relative to the case of a single walker across the entire energy domain. This plot is for Method 3 with a 25% energy sub-domain overlap. The error bars correspond to the standard deviation of time taken of 5 repeat WL simulations with different seeds on the PRNG. These data show that the optimal choices for efficiency are the 1 and 2 walker cases, with the 2 walker case marginally increasing the efficiency of each sub-domain. Additional walkers beyond this appear to confer no further benefit.
@@
Figure 4 shows a comparison of the speedup per WL instance (taken relative to one walker per energy sub-domain) for different numbers of walkers per energy sub-domain for Method 3, i.e. using statically-sized non-uniform energy sub-domains. It can be seen that useful speedup can be obtained, which is across all walker cases. In all cases, the efficiency remains super-linear for the majority if not for all sub-domains, however going beyond around 60 instance the efficiency drops below 100% while also appearing to plateau in terms of speedup. The initial improvement from the walkers can be attributed to the Central Limit Theorem [68]. By increasing the number of walkers, the number of samples increases (e.g. 2 walkers give twice as many samples, etc.) which leads to a reduction in statistical error on the estimate of the DOS proportional to $1/\sqrt{n}$, where $n$ is the number of samples. As the number of walkers increases this reduction in error goes as $1/\sqrt{2}$, $1/\sqrt{3}$, etc. which results in a quickly diminishing improvement in efficiency with respect to additional walkers. Our single sub-domain case with a varying quantity of independent walkers can be compared to the implementations found in Refs. [34, 35], however we only share statistics (i.e. the visit histogram and DOS) every 50 sweeps.
@@
Figure 5: Comparison of the walker efficiency per WL instance for different numbers of walkers per energy sub-domain, taken relative to the case of a single walker across the entire energy domain. This plot is for Method 3 with a 25% energy sub-domain overlap. The error bars correspond to the standard deviation of time taken of 5 repeat WL simulations with different seeds on the PRNG. The plot shows the efficiency of each walker introduced across 3 methods, with the 1 and 2 walker cases displaying the highest maximum efficiency per Wang–Landau instance.
@@
We then consider the efficiency of the walkers, as defined in Eq. 15 above. Figure 5 shows the efficiency of the instances seen in Figure 4. After the introduction of an additional walker, the maximum efficiency exceeds that of the single walker case, however going beyond two walkers per sub-domain, the efficiency begins to decrease substantially which would align with the previously outline Central Limit Theorem explanation. This increased efficiency from walkers stems partially from the minimum sub-domain size. Since the convergence time is limited by the slowest sub-domain, once the slowest sub-domain reaches the smallest allowed size, further addition of sub-domains will not result in increased efficiency. For the remainder of this study, all results presented are obtained using one walker per energy sub-domain.
III.3 Choice of method for construction of energy sub-domains
@@
Figure 6: Comparison of the energy sub-domain efficiency relative to the case of a single walker exploring the entire energy domain. Each method uses one walker per energy sub-domain and a 25% energy sub-domain overlap, and in all cases replica exchange was enabled. The error bars correspond to the standard deviation of time taken of 5 repeated WL simulations with different seeds on the PRNG. Panel (a) shows Method 1, i.e. dynamically load-balanced non-uniform sub-domains. Panel (b) shows Method 3, i.e. statically-sized non-uniform sub-domains. Panel (c) shows Method 5, i.e. uniformly-sized sub-domains. The plot shows the efficiency of each sub-domain introduced across 3 methods, with Method 1 showing the greatest and longest-lasting efficiency with its efficiency peak occurring at a smaller number of sub-domains. For the case of uniformly-sized energy sub-domains (Method 5), the sub-domain efficiency remains close to, or slightly below, unity regardless of the number of energy sub-domains used.
@@
Figure 7: A plot of the relative sizes of energy sub-domains for the proposed dynamic load-balancing scheme as a function of WL iteration number, to demonstrate the effect of load balancing on sub-domain size over the duration of a simulation. This plot is for Method 1 with one walker per energy sub-domain, a 25% energy sub-domain overlap, and a decomposition of the total energy domain into 6 sub-domains. The first iterations shows the initial uniform domain decomposition. It can be seen that the load balancing algorithm continues to make substantial adjustments to the relative sizes of energy sub-domains for the first few WL iterations, and that adjustment following an initial pre-sampling step is not enough to achieve good load balancing.
@@
Next, we consider the importance of the use of energy sub-domains. Figure 6 shows the sub-domain efficiency relative to a single walker operating within the entire energy domain. Within this plot we can see the sub-domain efficiency for Methods 1, 3 and 5 (from left to right respectively) across a range of sub-domains. When the three options for sizing of the energy sub-domains are considered, it can also be seen that the greatest increase in terms of performance is gained by implementing non-uniform energy domain decomposition. By further augmenting the sub-domains and introducing load balancing, a greater efficiency can be maintained at a higher $h$ as compared to using non-uniform energy domain decomposition with static sub-domain sizes. The performance gain observed from energy domain decomposition can be attributed to a reduction in the time taken for any walker to diffuse across the sub-domain, which reduces wasted sampling time spent on easily converged bins. Further discussion of this aspect can be found in Sec. III.4.
@@
To demonstrate the efficacy of our proposed dynamic energy sub-domain sizing scheme, we plot the comparative sizes of the energy sub-domains for an illustrative simulation run in Figure 7, for the case where a total of six energy sub-domains were used. The first iteration here corresponds to the pre-sampling step with initial uniform domain decomposition. It can be seen that there is a difference in the size of the sub-domains at iteration 2 as compared to iteration 13. This demonstrates the difficulty of a priori load balancing in the pre-sampling step and the benefits conferred by iterative load balancing. Also of note is the fact that the relative sizes of domain settle over the course of many iterations, demonstrating the numerical stability of our proposed load-balancing scheme.
III.4 Origins of observed super-linear efficiency
Now that greater than 100% efficiency for energy sub-domain decomposition has been demonstrated (evidenced in Figures 2 and 4), we support our initial analysis given in Sec. II.4 and confirm that super-linear efficiency results in a reduction in the total number of MC steps needed (summed over all WL instances) to converge the DOS. This is broadly equivalent to the statement that division of the total energy domain into sub-domains—and restricting walkers to operate in these sub-domains—reduces the total number of MC steps required to adequately explore the configuration space for the construction of the DOS.
Figure 8 shows the relative total number of MC steps required to complete all WL sampling iterations as a function of WL instance count, i.e. number of energy sub-domains, as compared to the case of a single walker operating across the entire energy domain. It can be seen that the total number of MC steps required to converge the simulation substantially drops once energy sub-domains are introduced. For the case of uniform energy sub-domains (panel (c)), as the number of sub-domains increases the total number of MC steps (sum of all steps taken across all $m\cdot h$ walkers) increases. In our conceptual model of a diffusive random walk in energy, this lack of reduction in MC steps required is due to the non-uniformity of ‘diffusivity’, $D$, quantifying the speed of walker exploration across the energy domain. In turn, this results in the WL convergence being dominated by the sub-domain which has the lowest $D$. However, for the two cases in which non-uniform energy sub-domains are employed (panels (a) and (b)), the total number of MC steps needed to fully converge the simulation decreases, which is particularly prominent for the first few additional WL instances, i.e. energy sub-domains. In this case, not only is the total work across the entire energy domain divided across WL instances, the total amount of work to be done is reduced (by around a factor of two for the case of two WL instances, i.e. energy sub-domains), which broadly agrees with the $\sim 4\times$ speedup seen in particular towards the left-hand data points of Figure 9. This analysis confirms our earlier assertion that the super-linear speedup, observed in both this study and in an earlier work, can be concretely understood in terms of a reduction in the total number of MC steps, i.e. the work, required to converge a given simulation.
@@ III.4 Origins of observed super-linear efficiency
Figure 8: MC steps taken to complete all WL sampling iterations relative to a single walker across the entire energy domain. Each method uses one walker and 25% overlap. Panel (a) shows results for Method 1, i.e. dynamically load-balanced non-uniform sub-domains. Panel (b) shows results for Method 3, i.e. statically-sized non-uniform sub-domains. Finally, panel (c) shows results for Method 5, i.e. uniformly-sized sub-domains. The plot shows that domain decomposition reduces the overall number of MC steps required to converge the WL sampling simulation.
III.5 Results summary
@@
Figure 9: Comparison of the speedup per WL instance, taken relative to the case of a single walker across the entire energy domain. The dotted line shows sub-domains squared $(h^{2})$. Each method has one walker and 25% overlap. The plot shows the greatest increase in performance can be obtained through non-uniform domain decomposition with load balancing further increasing performance.

### v3

@@ II.3 Model Implementation
As a test case for the parallelisation strategies studied in this work, we consider the application of WL sampling to the study of the phase behaviour of two selected metallic alloys: one well-studied binary system [54] and one multicomponent, high-entropy alloy (HEA). The application of WL to multicomponent alloys is an emerging area of research [55, 56, 57] and is therefore a relevant test case for this study, while the binary system provides a simple test case to compare against the HEA.
@@
In this work, we consider the application of WL sampling to a fixed lattice model describing the internal energy of a substitutional alloy. As benchmark systems, we select the AlTiCrMo high-entropy alloy, as well as the binary CuZn system. AlTiCrMo is classified as a refractory high-entropy superalloy [64] and forms a (partially) chemically ordered structure with decreasing temperature [65]. In the context of the present work, AlTiCrMo represents a complex system with competing interactions, multiple phase transitions, and a wide range of possible ordered configurations [66]. By contrast, CuZn is a well-studied binary alloy which unambiguously undergoes a single phase transition from a disordered bcc (A2) structure to an ordered B2 structure at intermediate temperatures [54]. It thus represents a simpler system than AlTiCrMo and allows us to examine how our proposed parallelisation strategies perform on model systems of varying complexity.
The parallelisation schemes considered in this study are implemented within the BraWl package [53], which focuses specifically on fixed-lattice simulations of substitutional alloys. In particular, the code implements the Bragg–Williams model [67, 68], which describes the internal energy of a given alloy as a sum over atom-atom effective pair interactions, which are assumed to have finite range. Explicitly, this Hamiltonian takes the form
@@
where $V_{i\alpha;j\alpha^{\prime}}$ denotes the effective pair interaction (EPI) between an atom of chemical species $\alpha$ on lattice site $i$ and an atom of chemical species $\alpha^{\prime}$ on lattice site $j$. $\{\xi_{i\alpha}\}$ indicates the site occupancies, where $\xi_{i\alpha}=1$ if site $i$ is occupied by an atom of chemical species $\alpha$, and $\xi_{i\alpha}=0$ otherwise. It is helpful to observe that this Hamiltonian can be thought of as a generalisation of the Lenz–Ising model used in elementary studies of magnetic phase transitions [69]. The EPIs of Eq. 12 are generally assumed to be homogeneous, isotropic, and of finite range, which simplifies evaluation of this expression. In this work, the EPIs used are obtained using the $S^{(2)}$ theory for multicomponent alloys, which is discussed extensively in earlier works [52, 70, 71]. For AlTiCrMo, we take the EPIs previously calculated in Ref. [66], which include interactions between pairs of atoms up to sixth-nearest neighbour distance and can be seen in panel (a) of Figure 2. For CuZn, we calculate EPIs at an optimised lattice constant of $a=2.971$ Å for the disordered bcc (A2) phase obtained using density functional theory (DFT) calculations with the generalised gradient approximation (GGA) to the exchange-correlation functional as parametrised by Perdew, Burke, and Ernzerhof [72]. The obtained EPIs can be seen in panel (b) of Figure 2. The Korringa–Kohn–Rostoker method is used [73], with the coherent potential approximation [74] employed to average over chemical disorder. In contrast to AlTiCrMo, for CuZn the EPIs are found to be much shorter in range and only significant on the first two coordination shells of the bcc lattice. We therefore truncate the EPIs for CuZn at this range.
@@ III Results and Discussion
In order to investigate the various outlined parallelisation schemes, the possible parallelisation schemes are grouped and labelled as ‘methods’, as shown in Table 1. Once grouped, the effect of the schemes on speedup per WL instance, efficiency per sub-domain and efficiency per walker, was investigated over a range of $h$ and $m$. The specific model systems being simulated were the equiatomic AlTiCrMo HEA, the phase behaviour of which has been investigated in an earlier work  [66], and the equiatomic CuZn alloy. The phase behaviour of these two systems is described in the Supplementary Material. We shall first begin by addressing the more complex of the two alloys, namely AlTiCrMo.
III.1 AlTiCrMo
In this benchmarking study assessing the efficiency of various parallelisation schemes, we use a simulation supercell made up of $4\times 4\times 4$ bcc cubic unit cells for a total of 128 atoms. We find that 128 atoms is sufficient to represent the sampling challenge with tests at larger lattice sizes ($6\times 6\times 6$ bcc cubic unit cells, 432 atoms) confirming the trends illustrated here. Our WL sampling simulations used $N=512$ energy bins over an energy domain of E=99E=-99 meV/atom to $E=0$ meV/atom (obtained via the method outlined in Sec. II.1 with $T_{\text{min}}=50K$ and $T_{\text{max}}=3000K$), completing 12 full WL iterations starting with $\log f_{1}=0.05$, where each iteration used half of the logarithm of the modification factor of the previous iteration. The AlTiCrMo HEA model is well-suited as a benchmark system as it exhibits two phase transitions which occur at two separate temperatures [66], and hence convergence of the DOS is non-trivial. We investigated performance for a range of parameters within our parallelisation schemes with the number of walkers per sub-domain $m$ varying from 1 to 6, the number of sub-domains $h$ varying from 1 to 16 and the sub-domain overlap varying from 0% to 75% for both overlap schemes. It is important to note that we impose a requirement that a minimum of two bins must lie in the overlap region. As such, the case which we refer to as 0% overlap still has two bins in the overlap region. This minimum overlap does not change based on the sub-domain size.
@@ III.1 AlTiCrMo
III.1.1 Effect of energy sub-domain overlap
Figure 3: Comparison of the speedup per WL instance for different choices of energy sub-domain overlap, taken relative to the case of a single walker across the entire energy domain for simulations of the AlTiCrMo test system. This plot is for simulations using Method 3 with a single walker per energy sub-domain. The dotted red line shows the square of the number of energy sub-domains, i.e., the number of WL instances used squared ($h^{2}$), while the solid red line shows the number of energy sub-domains, i.e., the number of WL instances ($h$). It can be seen that, for all choices of energy sub-domain overlap region size (excluding the 75% case), the speedup is significantly above 100% and that the choice of size of overlap region between energy sub-domains has noticeable impact on speedup.
@@ III.1.1 Effect of energy sub-domain overlap
Of the overlaps, the 25% to 50% overlap cases appear to be the ones with the most speedup, although not by a statistically significant amount. Such sub-domain overlap allows for better combining of the DOS, as well as a larger region for replica exchange, while still allowing walkers to focus more narrowly on their assigned sub-domain, resulting in more efficient sampling and higher speedup. The other cases introduce either too much or too little overlap, with too little overlap potentially harming the DOS combining and too much overlap leading to a greater portion of the configuration space being sampled by multiple walkers, which reduces the benefits of parallelisation. For the remainder of this study a 25% overlap is used. This deliberate choice allows for evaluation of performance with an overlap width which remains optimal to within the statistical uncertainty of our data, but still provides a significant range of overlap to realise any benefit from including replica exchange.
@@
It is important to note the super-linear scaling observed in some cases. This greater than 100% speedup per WL instance arises from the reduced number of total MC steps required when sampling due to the energy domain decomposition as explained in the previous section. This effect was discussed above, in Sec. II.4, and is also examined further in Sec. III.1.4, below. Of further note is the speedup for the first 3 WL instances matches the theoretical limit of $h^{2}$, derived in Sec. II.4 to within the uncertainty of our measurements.
@@
Figure 4: Normalised root mean square error (RMSE) as a function of the percentage overlap for simulations of the AlTiCrMo test system. The RMSE was obtained by comparing 23 iteration sampling runs with 16 windows to a well-converged 36 iteration sampling run with 1 window. The error bars correspond to the standard deviation of time taken of 5 repeat WL simulations. For each number of overlap bins, 5 separate 23 iteration simulations were performed and averaged. These data show that there is no correlation between the size of the energy sub-domain overlap region and the accuracy of the obtained global DOS up to a 50% overlap. Beyond 50% overlap the averaging power of the sampling is reduced hence the increase in normalised RMSE though not to a statistically significant extent.
@@
To assess the effect of overlap on DOS convergence, in Figure 4 we show the normalised root mean square error (RMSE) between a 36 iteration (well-converged) WL simulation in one sub-domain and 23 iteration WL simulations of varying overlap size with 16 sub-domains. It can be seen that there is no correlation between DOS convergence accuracy and the percentage overlap in the energy sub-domain overlap region up to 50%. Beyond this point the averaging power of the overlap regions is decreased resulting in an increased normalised RMSE. Of note, however, is that this increase in normalised RMSE increases from 0.15% to 0.3% which is not a statistically significant increase. While remaining at or below a 50% overlap, the primary benefit of the energy sub-domain overlap region(s) is to facilitate replica exchange and does not harm the sampling.
III.1.2 Influence of number of walkers used per energy sub-domain
@@
Figure 5: Comparison of the speedup per WL instance for different numbers of walkers per energy sub-domain, taken relative to the case of a single walker across the entire energy domain for simulations of the AlTiCrMo test system. This plot is for Method 3 with a 25% energy sub-domain overlap. The error bars correspond to the standard deviation of time taken of 5 repeat WL simulations with different seeds on the PRNG. These data show that the optimal choices for efficiency are the 1 and 2 walker cases, with the 2 walker case marginally increasing the efficiency of each sub-domain. Additional walkers beyond this appear to confer no further benefit.
@@
Figure 5 shows a comparison of the speedup per WL instance (taken relative to one walker per energy sub-domain) for different numbers of walkers per energy sub-domain for Method 3, i.e. using statically-sized non-uniform energy sub-domains. It can be seen that useful speedup can be obtained, which is across all walker cases. In all cases, the efficiency remains super-linear for the majority if not for all sub-domains, however going beyond around 60 instances the efficiency drops below 100% while also appearing to plateau in terms of speedup. The initial improvement from the walkers can be attributed to the Central Limit Theorem [76]. By increasing the number of walkers, the number of samples increases (e.g. 2 walkers give twice as many samples, etc.) which leads to a reduction in statistical error on the estimate of the DOS proportional to $1/\sqrt{n}$, where $n$ is the number of samples. As the number of walkers increases this reduction in error goes as $1/\sqrt{2}$, $1/\sqrt{3}$, etc. which results in a quickly diminishing improvement in efficiency with respect to additional walkers. Our single sub-domain case with a varying quantity of independent walkers can be compared to the implementations found in Refs. [35, 36], however we only share statistics (i.e. the visit histogram and DOS) every 50 sweeps.
@@
Figure 6: Comparison of the walker efficiency per WL instance for different numbers of walkers per energy sub-domain, taken relative to the case of a single walker across the entire energy domain for simulations of the AlTiCrMo test system. This plot is for Method 3 with a 25% energy sub-domain overlap. The error bars correspond to the standard deviation of time taken of 5 repeat WL simulations with different seeds on the PRNG. The plot shows the efficiency of each walker introduced across 3 methods, with the 1 and 2 walker cases displaying the highest maximum efficiency per Wang–Landau instance.
@@
We then consider the efficiency of the walkers, as defined in Eq. 15 above. Figure 6 shows the efficiency of the instances seen in Figure 5. After the introduction of an additional walker, the maximum efficiency exceeds that of the single walker case, however going beyond two walkers per sub-domain, the efficiency begins to decrease substantially which would align with the previously outlined Central Limit Theorem explanation. This increased efficiency from walkers stems partially from the minimum sub-domain size. Since the convergence time is limited by the slowest sub-domain, once the slowest sub-domain reaches the smallest allowed size, further addition of sub-domains will not result in increased efficiency. For the remainder of this study, all results presented are obtained using one walker per energy sub-domain.
III.1.3 Choice of method for construction of energy sub-domains
@@
Figure 7: Comparison of the energy sub-domain efficiency relative to the case of a single walker exploring the entire energy domain for simulations of the AlTiCrMo test system. Each method uses one walker per energy sub-domain and a 25% energy sub-domain overlap, and in all cases replica exchange was enabled. The error bars correspond to the standard deviation of time taken of 5 repeated WL simulations with different seeds on the PRNG. Panel (a) shows Method 1, i.e. dynamically load-balanced non-uniform sub-domains. Panel (b) shows Method 3, i.e. statically-sized non-uniform sub-domains. Panel (c) shows Method 5, i.e. uniformly-sized sub-domains. The plot shows the efficiency of each sub-domain introduced across 3 methods, with Method 1 showing the greatest and longest-lasting efficiency with its efficiency peak occurring at a smaller number of sub-domains. For the case of uniformly-sized energy sub-domains (Method 5), the sub-domain efficiency remains close to, or slightly below, unity regardless of the number of energy sub-domains used.
@@
Figure 8: A plot of the relative sizes of energy sub-domains for the proposed dynamic load-balancing scheme as a function of WL iteration number, to demonstrate the effect of load balancing on sub-domain size over the duration of a simulation for simulations of the AlTiCrMo test system. This plot is for Method 1 with one walker per energy sub-domain, a 25% energy sub-domain overlap, and a decomposition of the total energy domain into 6 sub-domains. The first few iterations show the initial uniform domain decomposition. It can be seen that the load balancing algorithm continues to make substantial adjustments to the relative sizes of energy sub-domains for the first few WL iterations, and that adjustment following an initial pre-sampling step is not enough to achieve good load balancing.
@@
Next, we consider the importance of the use of energy sub-domains. Figure 7 shows the sub-domain efficiency relative to a single walker operating within the entire energy domain. Within this plot we can see the sub-domain efficiency for Methods 1, 3 and 5 (from left to right respectively) across a range of sub-domains. When the three options for sizing of the energy sub-domains are considered, it can also be seen that the greatest increase in terms of performance is gained by implementing non-uniform energy domain decomposition. By further augmenting the sub-domains and introducing load balancing, a greater efficiency can be maintained at a higher $h$ as compared to using non-uniform energy domain decomposition with static sub-domain sizes. The performance gain observed from energy domain decomposition can be attributed to a reduction in the time taken for any walker to diffuse across the sub-domain, which reduces wasted sampling time spent on easily converged bins. Further discussion of this aspect can be found in Sec. III.1.4.
@@
To demonstrate the efficacy of our proposed dynamic energy sub-domain sizing scheme, we plot the comparative sizes of the energy sub-domains for an illustrative simulation run in Figure 8, for the case where a total of six energy sub-domains were used. The first iteration here corresponds to the pre-sampling step with initial uniform domain decomposition. It can be seen that there is a difference in the size of the sub-domains at iteration 2 as compared to iteration 13. This demonstrates the difficulty of a priori load balancing in the pre-sampling step and the benefits conferred by iterative load balancing. Also of note is the fact that the relative sizes of domains settle over the course of many iterations, demonstrating the numerical stability of our proposed load-balancing scheme.
@@
Figure 9: MC steps taken to complete all WL sampling iterations relative to a single walker across the entire energy domain for AlTiCrMo and CuZn. Each method uses one walker and 25% energy sub-domain overlap. Panel (i) shows results for Method 1, i.e. dynamically load-balanced non-uniform sub-domains. Panel (ii) shows results for Method 3, i.e. statically-sized non-uniform sub-domains. Finally, panel (iii) shows results for Method 5, i.e. uniformly-sized sub-domains. Both plots show that domain decomposition reduces the overall number of MC steps required to converge the WL sampling simulation.
@@
Figure 10: Comparison of the speedup per WL instance, taken relative to the case of a single walker across the entire energy domain, for simulations of AlTiCrMo and CuZn. The dotted line shows sub-domains squared $(h^{2})$. Each method has one walker and 25% overlap. Both plots show the greatest increase in performance can be obtained through non-uniform domain decomposition with a modest further increase from load balancing.
III.1.4 Origins of observed super-linear efficiency
Now that greater than 100% efficiency for energy sub-domain decomposition has been demonstrated (evidenced in Figures 3 and 5), we support our initial analysis given in Sec. II.4 and confirm that super-linear efficiency results in a reduction in the total number of MC steps needed (summed over all WL instances) to converge the DOS. This is broadly equivalent to the statement that division of the total energy domain into sub-domains—and restricting walkers to operate in these sub-domains—reduces the total number of MC steps required to adequately explore the configuration space for the construction of the DOS.
Figure 9 shows the relative total number of MC steps required to complete all WL sampling iterations as a function of WL instance count, i.e. number of energy sub-domains, as compared to the case of a single walker operating across the entire energy domain. For the AlTiCrMo case shown in panel (a), it can be seen that the total number of MC steps required to converge the simulation substantially drops once energy sub-domains are introduced. For the case of uniform energy sub-domains shown in panel (a.iii), as the number of sub-domains increases the total number of MC steps (sum of all steps taken across all $m\cdot h$ walkers) increases. In our conceptual model of a diffusive random walk in energy, this lack of reduction in MC steps required is due to the non-uniformity of ‘diffusivity’, $D$, quantifying the speed of walker exploration across the energy domain. In turn, this results in the WL convergence being dominated by the sub-domain which has the lowest $D$. However, for the two cases in which non-uniform energy sub-domains are employed (panels (a.i) and (a.ii)), the total number of MC steps needed to fully converge the simulation decreases, which is particularly prominent for the first few additional WL instances, i.e. energy sub-domains. In this case, not only is the total work across the entire energy domain divided across WL instances, the total amount of work to be done is reduced (by around a factor of two for the case of two WL instances, i.e. energy sub-domains), which broadly agrees with the $\sim 4\times$ speedup seen in particular towards the left-hand data points of panel (a) in Figure 10. This analysis confirms our earlier assertion that the super-linear speedup, observed in both this study and in an earlier work, can be concretely understood in terms of a reduction in the total number of MC steps, i.e. the work, required to converge a given simulation.
III.1.5 AlTiCrMo results summary